Purdue University MA460 Geometry Fall 2009 Syllabus Instructor: Uli Walther Office: 746 MATH Phone: 49-41959

                                E-mail: walther@math.purdue.edu

1 General Information

1.1 Class time / location:

TTh 10:30-11:45pm, 121 REC.

1.2 Office Hours:

Tue 2:30-3:30 and Th 1:30-2:30.

1.3 Textbook:

1. Course Notes on Course Webpage

2. Vol.1 of Euclid’s Elements (2nd Unabridged Edition, from Dover Publications)

1.4 Course Webpage:

Under

 http://www.math.purdue.edu/\~{}walther/teach/460/460-fall-tth.html

you find this syllabus as well as some other info concerning the course, including a link to our Project Rhea (see below).

1.5 Statement for Students with Disabilities

See the webpage, http://www.math.purdue.edu/~walther/teach.

1.6 Calculators

Calculators are meaningless in this course. They will not be used.

1.7 Course Grade:

Your course grade will be determined using the following distributions:

Rhea  5%
Quizzes 15%
HW  30%
Midterm 20%
Final  30%

The following are target gradelines: A ≥ 90%, B ≥ 80%, C ≥ 70%, D ≥ 60%.

2 Course Structure

If you disagree with the grading in any category, you need to resubmit the item in question with a written explanation stating why (and not just that) you deserve more credit.

2.1 Homework:

Homework will be collected every Thursday in class or in my office no later than 2:30pm. No late assignments will be accepted. Neither will homework deposited anywhere else.

Note, that homework cannot be done in one night. Start early.

In order to receive credit, homework must be readable, and stapled. Use one page per problem.

You are encouraged to attempt all the questions and discuss with your classmates. However, the write-up must be of your own.

On each assignement, some but not all problems will be graded, on a scale from 0-5 each. Ungraded solutions are worth one point.

2.2 Quizzes:

There will be a 20 minute quiz on many Thursdays. The first will be Thursday of week 2. The worst will be dropped.

2.3 Project Rhea

5% of the course grade comes from your contributions to Rhea (in the first 5 weeks. You are expected to make weekly contributions, each worth 1 percent. Contributions are due Thursday evening at midnight. For ideas what to contribute, see the Rhea page

https://kiwi.ecn.purdue.edu/rhea/index.php/MA460_(Fall2009Walther)

and watch for announcements in class.

2.4 Midterms:

There will be one midterm, Thursday, October 22 in class. If you should miss it, your final will carry extra weight.

2.5 Presentations:

Each week, some students may volunteer to present ungraded HW on the board. In the second half of the course, we’ll use groups to present some Euclid theorems. Each presentation is worth up to 0.5 % extra credit for HW and 1% for theorems.